# Towards Unified Latent Space for 3D Molecular Latent Diffusion Modeling

### v2

@@ 3.1 UAE-3D: Unified Variational Auto-Encoder for 3D Molecular Latent Diffusion Modeling
$\mathbf{Q}_{ij}=[\mathbf{H}^{n}_{i};\mathbf{H}^{e}_{ij}]\mathbf{W}^{q},$ $\mathbf{Q}\in\mathbb{R}^{|\mathcal{V}|\times|\mathcal{V}|\times d},$ (7)
$[\mathbf{K}_{ij};\mathbf{V}_{ij}]=[\mathbf{H}^{n}_{j};\mathbf{H}^{e}_{ij}]\mathbf{W}^{kv},$ $\mathbf{K},\mathbf{V}\in\mathbb{R}^{|\mathcal{V}|\times|\mathcal{V}|\times d},$ (8)
$\boldsymbol{\alpha}_{ij}=\text{softmax}_{j}\left(\frac{\mathbf{Q}_{ij}\mathbf{K}_{ij}^{\top}}{\sqrt{d}}\right),$ $\boldsymbol{\alpha}\in\mathbb{R}^{|\mathcal{V}|\times|\mathcal{V}|},$ (9)
$\hat{\mathbf{H}}_{i}^{n}=\sum_{j}\boldsymbol{\alpha}_{ij}\mathbf{V}_{ij},$ $\hat{\mathbf{H}}^{n}\in\mathbb{R}^{|\mathcal{V}|\times d},$ (10)
$\tilde{\mathbf{H}}_{i}^{n}=\operatorname{MLP}(\hat{\mathbf{H}}_{i}^{n}),$ $\tilde{\mathbf{H}}^{n}\in\mathbb{R}^{|\mathcal{V}|\times d},$ (11)
where $\mathbf{W}^{q}$ and $\mathbf{W}^{kv}$ are linear projectors. Unlike the original R-Trans, our implementation does not update the edge embedding $\mathbf{H}^{e}$ across layers. We find that this method is more efficient without compromising performance.
Compared to Transformer Vaswani et al. (2017), R-Trans can effectively integrate edge embeddings $\mathbf{H}^{e}\in\mathbb{R}^{|\mathcal{V}|\times|\mathcal{V}|\times d}$ and atom embeddings $\mathbf{H}^{n}\in\mathbb{R}^{|\mathcal{V}|\times d}$, despite their different shapes, making it well-suited for 3D molecule encoding. This integration occurs during the computation of queries, keys and values (Equation (7) and (8)), ensuring the output atom embeddings fully incorporate edge information.
Decoder. We employ a standard Transformer Vaswani et al. (2017) as decoder, with three additional predictors to reconstruct complete 3D molecules: atom types, bond types, and atom coordinates. Unlike the encoder, our decoder includes no molecule-specialized design, because it processes sequences of latents. Given the encoder’s latents $\mathbf{Z}$, our Transformer decoder provides the representation:
$\mathbf{P}=\operatorname{Transformer}(\mathbf{Z})\in\mathbb{R}^{|\mathcal{V}|\times d}.$ (12)
We then obtain predictions for three molecular modalities:
$\hat{\mathbf{X}}_{i}$ $=\operatorname{MLP}_{1}(\mathbf{P}_{i})$ $\in\mathbb{R}^{3},$ (13)
$\hat{\mathbf{F}}_{i}$ $=\operatorname{MLP}_{2}(\mathbf{P}_{i})$ $\in\mathbb{R}^{N_{a}},$ (14)
$\hat{\mathbf{E}}_{ij}$ $=\operatorname{MLP}_{3}(\mathbf{P}_{i}+\mathbf{P}_{j})$ $\in\mathbb{R}^{N_{b}},$ (15)
@@
Reconstruction Loss. We define the distance between our reconstructed graph and the original graph $|\hat{\mathbf{G}}-\mathbf{G}|$, which includes multiple components to ensure the fidelity of the generated 3D molecules:
$\mathcal{L}_{\text{atom}}$ $=\frac{1}{|\mathcal{V}|}\sum_{i\in\mathcal{V}}\operatorname{CrossEntropy}(\hat{\mathbf{F}}_{i},\mathbf{F}_{i}),$ (16)
$\mathcal{L}_{\text{coordinate}}$ $=\frac{1}{|\mathcal{V}|}\sum_{i\in\mathcal{V}}\operatorname{MSE}(\hat{\mathbf{X}}_{i},\mathbf{X}_{i}),$ (17)
$\mathcal{L}_{\text{bond}}$ $=\frac{1}{|\mathcal{V}|^{2}}\sum_{i,j\in\mathcal{V}}\operatorname{CrossEntropy}(\hat{\mathbf{E}}_{ij},\mathbf{E}_{ij}),$ (18)
$\mathcal{L}_{\text{distance}}$ $=\frac{1}{|\mathcal{V}|^{2}}\sum_{i,j\in\mathcal{V}}\operatorname{MSE}(\|\hat{\mathbf{X}}_{i}-\hat{\mathbf{X}}_{j}\|,\mathbf{D}_{ij}),$ (19)
where $\mathcal{L}_{\text{atom}}$ and $\mathcal{L}_{\text{bond}}$ are cross-entropy losses for atom and bond types; $\mathcal{L}_{\text{coordinate}}$ measures the MSE between the predicted and the ground truth coordinates; and $\mathcal{L}_{\text{distance}}$ serves as an additional constraint on the predicted coordinates, to ensure correct inter-atomic distances. In particular, we introduce a bonded distance loss which is incorporated within $\mathcal{L}_{\text{distance}}$. In our implementation, we assign a higher weight to the loss computed over bonded atom pairs to enforce stricter geometric constraints where chemical bonds exist. This weighted scheme ensures that the model captures the intrinsic molecular geometry more accurately and improves the overall VAE reconstruction quality. The overall reconstruction loss is a weighted sum of these components:
$\mathcal{L}_{\text{recon}}=\boldsymbol{\gamma}\cdot[\mathcal{L}_{\text{atom}},\mathcal{L}_{\text{bond}},\mathcal{L}_{\text{coordinate}},\mathcal{L}_{\text{distance}}]^{\top},$ (20)
@@
$\mathcal{L}_{\text{UAE-3D}}=$ $\mathcal{L}_{\text{recon}}$ (21)
$+$ $\beta\cdot\mathbb{E}_{\mathbf{G}\sim p_{\text{data}}}\left[D_{\operatorname{KL}}(q(\mathbf{Z}|\mathbf{G})\|p(\mathbf{Z}))\right],$
@@
SE(3)-Equivariant Augmentations. To ensure the VAE follows SE(3)-equivariance, we apply $\text{SE}(3)$ transformations $\mathbf{R}\in\text{SE}(3)$ as data augmentation on the molecular 3D coordinates $\mathbf{X}$ Abramson et al. (2024). The VAE $\mathcal{D}(\mathcal{E}(\cdot))$ is trained to be equivariant to these transformations:
$\mathcal{L}_{\text{recon}}=\|\mathcal{D}(\mathcal{E}(\mathbf{R}\circ\mathbf{G}))-\mathbf{R}\circ\mathbf{G}\|,$ (22)
where $\mathbf{R}\circ\mathbf{G}=\langle\mathbf{F},\mathbf{E},\mathbf{R}(\mathbf{X})\rangle$ denotes applying the transformation $\mathbf{R}$ on the 3D coordinates of $\mathbf{G}$. In practice, $\mathbf{R}$ combines a random rotation from $\text{SO}(3)$ and a random translation from $\mathcal{N}(\mathbf{0},0.01\mathbf{I}_{3})$, where $0.01$ controls the range of the translation augmentation. When these SE(3) augmentations are applied, Equation (22) becomes equivalent to Equation (20), training the VAE to follow SE(3)-equivariance.
3.2 UDM-3D: Unified Latent Diffusion Modeling for 3D Molecule Generation
@@
Diffusion Transformer (DiT). Given the unified latent space provided by our UAE-3D model, we now apply the DiT Peebles and Xie (2023) as our diffusion model $\boldsymbol{\epsilon}_{\theta}$ for 3D molecular latent generation. DiT’s promising performance in image generation underscores its superior capability in modeling latent sequences. Unlike a standard Transformer Vaswani et al. (2017), DiT replaces the layernorm Ba et al. (2016) by adaptive layernorm (adaLN) Perez et al. (2018), where the shift $\mathbf{y}$ and scale $\mathbf{b}$ parameters are dynamically generated based on the diffusion timestep $t$: $[\mathbf{y};\mathbf{b}]=\text{MLP}(\text{Embed}_{t}(t))$, where $\text{Embed}(\cdot)$ embeds timestep $t$ into a vector. These shift $\mathbf{y}$ and scale $\mathbf{b}$ parameters act as “gates”, selectively activating different dimension given different timesteps, enabling effective dynamic-scale denoising in diffusion.
Conditional Generation with Classifier-Free Guidance. For conditional generation, we directly combine the condition embedding $\text{Embed}_{c}(\mathbf{c})$ with the time embedding $\text{Embed}_{t}(\mathbf{t})$ when generating parameters for the adaLN:
$[\mathbf{y};\mathbf{b}]=\text{MLP}(\text{Embed}_{t}(t)+\text{Embed}_{c}(c)).$ (23)
In this way, we effectively inform the diffusion model $\boldsymbol{\epsilon}_{\theta}(\mathbf{Z}^{(t)},\mathbf{t},\mathbf{c})$ with the condition information.
@@
To further enforce conditioning during inference, we employ the classifier-free guidance Ho and Salimans (2022) (CFG) to find $\mathbf{Z}$ that has a high $\log p(\mathbf{c}|\mathbf{Z})$. By Bayes’ rule, the gradient of this objective is $\nabla_{\mathbf{Z}}\log p(\mathbf{c}|\mathbf{Z})\propto\nabla_{\mathbf{Z}}\log p(\mathbf{Z}|\mathbf{c})-\nabla_{\mathbf{Z}}\log p(\mathbf{Z})$. Following Song et al. (2021), we can interprete the diffusion model $\boldsymbol{\epsilon}_{\theta}$’s output as score functions, and therefore maximize $p(\mathbf{c}|\mathbf{Z})$ using the modified denoising function:
$\tilde{\boldsymbol{\epsilon}}_{\theta}(\mathbf{Z}^{(t)},\mathbf{t},\mathbf{c})=(1+w)\boldsymbol{\epsilon}_{\theta}(\mathbf{Z}^{(t)},\mathbf{t},\mathbf{c})-w\boldsymbol{\epsilon}_{\theta}(\mathbf{Z}^{(t)},\mathbf{t}),$ (24)
@@
where $w\in[0,+\infty)$ is a hyperparameter controlling the guidance strength; $\boldsymbol{\epsilon}_{\theta}(\mathbf{Z}^{(t)},\mathbf{t},\mathbf{c})$ is the conditional variant of $\boldsymbol{\epsilon}_{\theta}$, incorporating the property $\mathbf{c}$; $\boldsymbol{\epsilon}_{\theta}(\mathbf{Z}^{(t)},\mathbf{t})$ is the unconditioned variant that ignores $\mathbf{c}$. To train $\boldsymbol{\epsilon}_{\theta}$, we randomly drop condition $\mathbf{c}$ with a certain probability, allowing $\boldsymbol{\epsilon}_{\theta}$ to learn both conditional and unconditional distributions.
@@
Table 2: Performance of de novo 3D molecule generation on GEOM-Drugs. * indicates results reproduced using official source codes, while other baseline results are taken from Huang et al. (2023b).
@@
Train 1.000 1.000 1.000 1.000 0.000 0.585 0.999 0.584 00.251
CDGS 0.991 0.706 0.285 0.285 0.285 0.262 0.789 0.022 22.051
JODO 1.000 0.981 0.874 0.905 0.902 0.417 0.993 0.483 02.523
MiDi* 0.968 0.822 0.633 0.654 0.652 0.392 0.951 0.196 07.054
EQGAT-diff* 1.000 0.989 0.845 0.863 0.859 0.377 0.983 0.161 05.898
UDM-3D, ours 1.000 0.925 0.879 0.913 0.907 0.525 0.990 0.540 00.692
@@
Train 0.861 0.028 13.73 1.56E-04 1.81E-04 1.56E-04
EDM 0.831 0.002 31.29 4.29E-01 4.96E-01 1.46E-02
JODO 0.845 0.010 19.99 8.49E-02 1.15E-02 6.68E-04
MiDi* 0.750 0.003 23.14 1.17E-01 9.57E-02 4.46E-03
GeoLDM 0.843 0.008 30.68 3.91E-01 4.22E-01 1.69E-02
EQGAT-diff* 0.825 0.007 26.33 1.55E-01 5.21E-02 2.10E-03
UDM-3D, ours 0.852 0.014 17.36 9.89E-03 5.11E-03 1.74E-04
Table 3: Performance of de novo 3D molecule generation on QM9. * indicates results reproduced using official source codes, while other baseline results are taken from Huang et al. (2023b).
@@
Train 0.999 0.988 0.989 0.989 0.000 0.490 0.992 0.946 0.063
CDGS 0.997 0.951 0.951 0.936 0.860* 0.493 0.973 0.784 0.798
JODO 0.999 0.988 0.990 0.960 0.780* 0.522 0.986 0.934 0.138
MiDi* 0.998 0.976 0.980 0.954 0.769 0.501 0.979 0.882 0.187
EQGAT-diff* 0.999 0.971 0.965 0.950 0.891 0.482 0.950 0.703 2.088
UDM-3D, ours 0.999 0.988 0.983 0.972 0.948 0.508 0.987 0.897 0.161
[…]
Train 0.994 0.953 0.877 5.44E-04 4.65E-04 1.78E-04
EDM 0.986 0.817 1.285 1.30E-01 1.82E-02 6.64E-04
MDM 0.992 0.896 4.861 2.74E-01 6.60E-02 2.39E-02
JODO 0.992 0.934 0.885 1.48E-01 1.21E-02 6.29E-04
GeoLDM 0.989 0.897 1.030 2.40E-01 1.00E-02 6.59E-04
MiDi* 0.983 0.842 1.100 8.96E-01 2.08E-02 8.14E-04
EQGAT-diff* 0.988 0.888 1.520 4.21E-01 1.89E-02 1.24E-03
UDM-3D, ours 0.993 0.929 0.881 7.04E-02 9.84E-03 3.47E-04
Table 4: Mean Absolute Error (MAE) for conditional 3D molecule generation on QM9. Lower values represent better performance.
U-Bound 1.613 8.98 6.879 645 1457 1464
EDM 1.123 2.78 1.065 371 0601 0671
EEGSDE 0.777 2.50 0.941 302 0447 0487
GeoLDM 1.108 2.37 1.025 340 0522 0587
GeoBFN 0.998 2.34 0.949 328 0516 0577
JODO 0.628 1.42 0.581 226 0256 0335
UDM-3D, ours 0.603 1.76 0.553 216 0247 0313
L-Bound 0.043 0.09 0.040 039 0036 0065

### v3

@@ 3.1 Unified Variational Auto-Encoder for 3D Molecular Latent Diffusion Modeling
$\mathbf{Q}_{ij}=[\mathbf{H}^{n}_{i};\mathbf{H}^{e}_{ij}]\mathbf{W}^{q}\in\mathbb{R}^{d},$ (5) $[\mathbf{K}_{ij};\mathbf{V}_{ij}]=[\mathbf{H}^{n}_{j};\mathbf{H}^{e}_{ij}]\mathbf{W}^{kv}\in\mathbb{R}^{2d},$ (6)
$\boldsymbol{\alpha}_{ij}=\text{softmax}_{j}\left(\frac{\mathbf{Q}_{ij}\mathbf{K}_{ij}^{\top}}{\sqrt{d}}\right)\in\mathbb{R},$ (7) $\hat{\mathbf{H}}_{i}^{n}=\operatorname{MLP}(\sum_{j}\boldsymbol{\alpha}_{ij}\mathbf{V}_{ij})\in\mathbb{R}^{d},$ (8)
where $\mathbf{W}^{q}$ and $\mathbf{W}^{kv}$ are linear projectors. Unlike the original R-Trans, our implementation does not update the edge $\mathbf{H}^{e}$ every layer, which improves efficiency without compromising performance.
Compared to Transformer [22], R-Trans is well-suited for 3D molecule encoding, that it can effectively integrate edge embeddings $\mathbf{H}^{e}\in\mathbb{R}^{|\mathcal{V}|\times|\mathcal{V}|\times d}$ and atom embeddings $\mathbf{H}^{n}\in\mathbb{R}^{|\mathcal{V}|\times d}$, despite their different shapes. This integration occurs during the computation of queries, keys and values (Equation (5) and (6)), ensuring the output atom embeddings fully incorporate edge information.
Decoder. Given encoder’s latents $\mathbf{Z}$, we employ a Transformer [22] decoder $\mathcal{D}(\cdot)$ to obtain the atom-wise output $\mathbf{P}=\operatorname{Transformer}(\mathbf{Z})\in\mathbb{R}^{|\mathcal{V}|\times d}$. Unlike the encoder, our decoder includes no molecule-specialized design, because it processes sequences of latents. Further, we employ three MLP predictors for atom, bond types, and atom coordinates to reconstruct complete 3D molecules:
$\hat{\mathbf{X}}_{i}=\operatorname{MLP}_{1}(\mathbf{P}_{i})\in\mathbb{R}^{3};\quad\hat{\mathbf{F}}_{i}=\operatorname{MLP}_{2}(\mathbf{P}_{i})\in\mathbb{R}^{N_{a}};\quad\hat{\mathbf{E}}_{ij}=\operatorname{MLP}_{3}(\mathbf{P}_{i}+\mathbf{P}_{j})\in\mathbb{R}^{N_{b}},$ (9)
@@
Reconstruction Loss. We define the distance between our reconstructed graph and the original graph $\|\hat{\mathbf{G}}-\mathbf{G}\|$. It includes multiple components to ensure the fidelity of generated 3D molecules:
$\mathcal{L}_{\text{atom}}$ $=\frac{1}{|\mathcal{V}|}\sum_{i\in\mathcal{V}}\operatorname{CE}(\hat{\mathbf{F}}_{i},\mathbf{F}_{i}),$ $\mathcal{L}_{\text{coordinate}}$ $=\frac{1}{|\mathcal{V}|}\sum_{i\in\mathcal{V}}\operatorname{MSE}(\hat{\mathbf{X}}_{i},\mathbf{X}_{i}),$ (10)
$\mathcal{L}_{\text{bond}}$ $=\frac{1}{|\mathcal{V}|^{2}}\sum_{i,j\in\mathcal{V}}\operatorname{CE}(\hat{\mathbf{E}}_{ij},\mathbf{E}_{ij}),$ $\mathcal{L}_{\text{distance}}$ $=\frac{1}{|\mathcal{V}|^{2}}\sum_{i,j\in\mathcal{V}}w_{ij}\operatorname{MSE}(\|\hat{\mathbf{X}}_{i}-\hat{\mathbf{X}}_{j}\|,\mathbf{D}_{ij}).$
where $\operatorname{CE}$ denotes cross-entropy loss, and $\mathcal{L}_{\text{atom}}$ and $\mathcal{L}_{\text{bond}}$ are losses for atom and bond types; $\mathcal{L}_{\text{coordinate}}$ measures the MSE between the predicted and the ground truth coordinates; and $\mathcal{L}_{\text{distance}}$ serves as an extra constraint on coordinate predictions, to ensure correct inter-atomic distances. Notably, we incorporate a bonded distance loss inside $\mathcal{L}_{\text{distance}}$ by adjusting the pairwise weight $w_{ij}$:
$w_{ij}=\begin{cases}1+\lambda,&\text{if atom }i\text{ and }j\text{ are bonded,}\\ 1,&\text{otherwise.}\end{cases}$ (11)
This design prioritizes accurate inter-atomic distances between bonded atoms, enforcing stricter geometry in chemically important regions. $\lambda$ is a hyperparameter that controls this constraint’s strength. Given the components above, we define the distance between our reconstructed graph and the original graph (i.e., the reconstruction loss) as follows:
$\mathcal{L}_{\text{recon}}=\mathbb{E}_{\mathbf{G}\sim p_{\text{data}}}\|\hat{\mathbf{G}}-\mathbf{G}\|=\mathbb{E}_{\mathbf{G}\sim p_{\text{data}}}\boldsymbol{\gamma}\cdot[\mathcal{L}_{\text{atom}},\mathcal{L}_{\text{bond}},\mathcal{L}_{\text{coordinate}},\mathcal{L}_{\text{distance}}]^{\top},$ (12)
@@
$\mathcal{L}_{\text{UAE-3D}}=\mathcal{L}_{\text{recon}}\\ +\beta\cdot\mathbb{E}_{\mathbf{G}\sim p_{\text{data}}}\left[D_{\operatorname{KL}}(q(\mathbf{Z}|\mathbf{G})\|p(\mathbf{Z}))\right],$ (13)
@@
SE(3)-Equivariant Augmentations. To enforce SE(3)-equivariance in UAE-3D, we apply random transformations $\mathbf{R}\in\text{SE}(3)$ to input coordinates $\mathbf{X}$ during training. Each $\mathbf{R}$ combines a rotation from $\text{SO}(3)$ and a random translation from $\mathcal{N}(\mathbf{0},0.01\mathbf{I}_{3})$. Crucially, the reconstruction loss $\mathcal{L}_{\text{recon}}$ (cf. Equation (12)) remains unchanged but operates on transformed inputs $\mathbf{R}\circ\mathbf{G}=\langle\mathbf{F},\mathbf{E},\mathbf{R}(\mathbf{X})\rangle$. This process can be informally conceptualized as a reconstruction loss of $\|\mathcal{D}(\mathcal{E}(\mathbf{R}\circ\mathbf{G}))-\mathbf{R}\circ\mathbf{G}\|$, where the autoencoder $\mathcal{D}(\mathcal{E}(\cdot))$ learns to preserve geometric consistency under SE(3) transformations.
3.2 Unified Latent Diffusion Modeling for 3D Molecule Generation
@@
Diffusion Transformer (DiT). Given the unified latent space provided by our UAE-3D model, we adopt the DiT [23] as the backbone diffusion model $\boldsymbol{\epsilon}_{\theta}$ for 3D molecular latent generation. Originally developed for image LDM, DiT has demonstrated strong performance in modeling latent sequences. Compared to a standard Transformer [22], DiT replaces the layernorm [28] by adaptive layernorm (adaLN) [29], where the scale $\mathbf{y}\in\mathbb{R}^{d}$ and shift $\mathbf{b}\in\mathbb{R}^{d}$ parameters are conditioned on the diffusion timestep $t\in(0,1]$. Specifically, $[\mathbf{y};\mathbf{b}]=\text{MLP}_{\text{adaLN}}(\text{Embed}_{t}(t))$, where $\text{Embed}_{t}(\cdot)$ is a shared linear layer and $\text{MLP}_{\text{adaLN}}$ is a module specific to each adaLN instance. These generated shift and scale parameters act as “soft gates”, enabling timestep-dependent activation of DiT’s hidden representations and facilitating more effective, scale-adaptive denoising.
Conditional Generation with Classifier-Free Guidance. To enable generation with a condition vector $\mathbf{c}$, we extend the adaLN modules by combining the condition embedding $\text{Embed}_{c}(\mathbf{c})$ with the timestep embedding $\text{Embed}_{t}(t)$ when computing the scale $\mathbf{y}$ and shift $\mathbf{b}$ parameters. This allows the diffusion model $\boldsymbol{\epsilon}\theta(\mathbf{Z}^{(t)},t,\mathbf{c})$ to incorporate conditioning information during denoising:
$[\mathbf{y};\mathbf{b}]=\text{MLP}_{\text{adaLN}}(\text{Embed}_{t}(t)+\text{Embed}_{c}(\mathbf{c})).$ (14)
@@
To further enforce conditioning during inference, we employ the classifier-free guidance (CFG) [30] to find $\mathbf{Z}$ that has a high $\log p(\mathbf{c}|\mathbf{Z})$. By Bayes’ rule, the gradient of this objective is $\nabla_{\mathbf{Z}}\log p(\mathbf{c}|\mathbf{Z})\propto\nabla_{\mathbf{Z}}\log p(\mathbf{Z}|\mathbf{c})-\nabla_{\mathbf{Z}}\log p(\mathbf{Z})$. Following [31], we can interprete the diffusion model $\boldsymbol{\epsilon}_{\theta}$’s output as score functions, and therefore maximize $p(\mathbf{c}|\mathbf{Z})$ using the modified denoising function:
$\tilde{\boldsymbol{\epsilon}}_{\theta}(\mathbf{Z}^{(t)},t,\mathbf{c})=(1+w)\boldsymbol{\epsilon}_{\theta}(\mathbf{Z}^{(t)},t,\mathbf{c})-w\boldsymbol{\epsilon}_{\theta}(\mathbf{Z}^{(t)},t),$ (15)
@@
where $w\in[0,+\infty)$ is a hyperparameter controlling the guidance strength; $\boldsymbol{\epsilon}_{\theta}(\mathbf{Z}^{(t)},\mathbf{t},\mathbf{c})$ is the conditional variant of $\boldsymbol{\epsilon}_{\theta}$, incorporating the property $\mathbf{c}$; $\boldsymbol{\epsilon}_{\theta}(\mathbf{Z}^{(t)},\mathbf{t})$ is the unconditioned variant that ignores $\mathbf{c}$. To train $\boldsymbol{\epsilon}_{\theta}$, we randomly drop condition $\mathbf{c}$ with a certain probability $p_{\text{drop}}$, allowing $\boldsymbol{\epsilon}_{\theta}$ to learn both conditional and unconditional distributions.
@@
Table 1: Performance of de novo 3D molecule generation on GEOM-Drugs. * indicates results reproduced using official source codes, while other baseline results are taken from [6].
@@
Train 00.251 1.000 1.000 1.000 1.000 0.000 0.585 0.999 0.584
CDGS 22.051 0.991 0.706 0.285 0.285 0.285 0.262 0.789 0.022
JODO 2.523 1.000 0.981 0.874 0.905 0.902 0.417 0.993 0.483
MiDi* 7.054 0.968 0.822 0.633 0.654 0.652 0.392 0.951 0.196
EQGAT-diff* 5.898 1.000 0.989 0.845 0.863 0.859 0.377 0.983 0.161
@@
${}_{\text{3D}}$$\downarrow$ AtomStable MolStable Bond $\downarrow$ Bond $\downarrow$ Dihedral $\downarrow$
Train 13.73 0.861 0.028 1.56E-04 1.81E-04 1.56E-04
EDM 31.29 0.831 0.002 4.29E-01 4.96E-01 1.46E-02
JODO 19.99 0.845 0.010 8.49E-02 1.15E-02 6.68E-04
MiDi* 23.14 0.750 0.003 1.17E-01 9.57E-02 4.46E-03
GeoLDM 30.68 0.843 0.008 3.91E-01 4.22E-01 1.69E-02
EQGAT-diff* 26.33 0.825 0.007 1.55E-01 5.21E-02 2.10E-03